Advertisements
Advertisements
प्रश्न
In a lottery ticket, there are 20 prizes and 25 blanks.
Statement (1): Probability of not getting the prize \[=1-\frac{20}{45}\].
Statement (2): P(getting prize) + P(blank) = 1
विकल्प
Both the statements are true.
Both the statements are false.
Statement 1 is true, and statement 2 is false.
Statement 1 is false, and statement 2 is true.
Advertisements
उत्तर
Both the statements are true.
Explanation:
Given, tickets = 20 prizes and 25 blanks
total tickets = 20 + 25 = 45
By formula; the probability of an event = \[\frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}}\]
P(getting prize) = \[\frac{20}{45}\]
P(blank) = \[\frac{25}{45}\]
P(getting prize) + P(blank) = \[\frac{20}{45} + \frac{25}{45}\]
\[{} = \frac{20 + 25}{45}\]
\[{} = \frac{45}{45}\]
\[{} = 1.\]
Thus, P(getting prize) + P(blank) = 1
So, statement 2 is true.
⇒ P(blank) = 1 − P(getting prize)
⇒ P(blank) = \[1 - \frac{20}{45}\]
We know that,
P(not getting the prize) = P(blank)
So, statement 1 is true.
∴ Both the statements are true.
